# Menger and consonant sets in the Sacks model

### v2

@@ 6. The weak grouped Menger game and Sacks forcing
For the proof of Theorem 6.1, we need the following auxiliary result. Similarly to Lemma 4.2, it can be proved in the same way as [7, Lemma 5.10], a rather standard argument is left to the reader.
@@
(6.2.2) $X=\bigcup\big{\{}L\subseteq 2^{\omega}\>:\>L\subseteq X,\mbox{ $L$ is compact,and $L$ is coded in }V[G_{\alpha}]\big{\}}\mbox{ and}$
(6.2.3) $2^{\omega}\setminus X=\bigcup\big{\{}L\subseteq 2^{\omega}\>:\>L\subseteq 2^{\omega}\setminus X,\mbox{ $L$ is compact, and $L$ is coded in }V[G_{\alpha}]\big{\}}.$
@@
Take $p$ and $F_{n},k_{n},l_{n},y_{\sigma_{n}}$ from Lemma 3.7, applied to $p_{0}$ and $\tau$. Let $\Sigma_{n}$ be the set of all maps $\sigma\colon F_{n}\to 2^{k_{n}}$ consistent with $p$, where $n\in\omega$. By Lemma 3.9(1), for each map $\sigma\in\Sigma_{n}$, there are maps $\sigma^{\prime}\neq\sigma^{\prime\prime}\in\Sigma_{n+1}$ extending $\sigma$, which implies $[\sigma^{\prime}]\cap[\sigma^{\prime\prime}]=\emptyset$. For $S:=\operatorname{supp}(p)$, define a perfect set $K$, exactly in the same way as before Lemma 2.1, i.e.,
@@
Then the family of all sets $[\sigma]$, where $\sigma\in\bigcup_{n\in\omega}\Sigma_{n}$, is a basis for $K$. Note also that Lemma 3.9(3) ensures that $S$ and the sequence ${\langle\,k_{n},F_{n},\Sigma_{n}:n\in\omega\,\rangle}$ satisfy $(e_{f})$ stated at the beginning of Section 2.
@@
Again, take $p$ and $F_{n},k_{n},l_{n},y_{\sigma_{n}}$ from Lemma 3.7, applied to $p_{0}$ and $\tau$. Let $\Sigma_{n}$, $[\sigma]$ for $\sigma\in\Sigma_{n}$, $K$, and $h:K\to 2^{\omega}$ be defined in the same way as in the proof of (6.2.2). We have $p\Vdash\tau\in h[K].$
@@
The first half of Corollary 6.3(2), namely that all perfectly meager subsets of ${2^{\omega}}$ have cardinality at most $\omega_{1}$ in the Sacks model, was established by Miller [17, § 5].
@@
In $V[G_{\omega_{2}}]$, considered in the above corollary, there is a Luzin subset of ${2^{\omega}}$, i.e., an uncountable set whose intersection with any meager set is at most countable, which is totally imperfect and Menger [14], but it is not (perfectly) meager. There is also a perfectly meager set that is not Menger: Since $\mathfrak{d}=\omega_{1}$ in $V[G_{\omega_{2}}]$, there is a dominating set $X=\{\,x_{\alpha}:\alpha<\omega_{1}\,\}$ in ${[\omega]^{\omega}}$, where $x_{\beta}\leq^{*}x_{\alpha}$ for all ordinal numbers $\beta<\alpha<\omega_{1}$. This set is not Menger. The set $X\cup\mathrm{Fin}$ satisfies the Hurewicz covering property [3, Theorem 10.], and any totally imperfect set with this property is perfectly meager [16, Theorem 5.5.]. Thus, the set $X$ is perfectly meager, too. Alternatively, we could use here the main result of [20], which implies directly that $X$ is perfectly meager since $\leq^{*}$ is a Borel subset of $\omega^{\omega}\times\omega^{\omega}$.
@@ 7. Menger sets and Hechler forcing
The results from the previous section lead to the question, whether $\mathfrak{d}<\mathfrak{c}$ implies that any totally imperfect Menger subset of ${2^{\omega}}$ has cardinality at most $\mathfrak{d}$. We address this problem, showing that this is not the case. We also provide a consistent result that the size of the family of all Hurewicz subsets of ${2^{\omega}}$ can be equal $2^{\mathfrak{c}}$ even if $\mathfrak{d}<\mathfrak{c}$. Let $\mathbb{P}$ be a definable ccc forcing notion of size $\mathfrak{c}$ which adds dominating reals over a ground model (e.g., the poset defined in [4, p. 95], nowadays commonly named Hechler forcing) and $\mathbb{P}_{\omega_{1}}$ be an iterated forcing of length $\omega_{1}$ with finite support, where each iterand is equal to $\mathbb{P}$. In this section, by $G_{\omega_{1}}$ we mean a $\mathbb{P}_{\omega_{1}}$-generic filter over a ground model $V$ of ZFC.
@@
Let $X\subseteq{2^{\omega}}\cap V$ and $\phi\colon X\to{[\omega]^{\omega}}$ be a continuous function. The function $\phi$ can be extended to a continuous function $\Phi$ defined on a $G_{\delta}$-set $A$, containing $X$. Then there is an ordinal number $\alpha<\omega_{1}$ such that the function $\Phi$ and the set $A$ are coded in $V[G_{\alpha}]$. It follows that $\Phi[X]\subseteq V[G_{\alpha}]$. Since there exists a function $g_{\alpha}$, added in step $\alpha$, which dominates any real from $V[G_{\alpha}]$, the set $\Phi[X]$ is bounded in $V[G_{\omega_{1}}]$. By the Hurewicz–Recław characterization of the Hurewicz property [21, Proposition 1], the set $X$ is Hurewicz [^5] As noted by the referee, a similar argument actually gives that $\phi[X]$ is bounded for any Borel function $f:X\to\omega^{\omega}$. in $V[G_{\omega_{1}}]$. ∎
@@
Assume that $V$ satisfies $\neg$¬CH. In $V[G_{\omega_{1}}]$, we have $\mathfrak{d}<\mathfrak{c}$ and the following assertions hold.
@@
Since Cohen reals are added by any tail of the considered iteration, any subset of $2^{\omega}\cap V$ is also Rothberger in $V[G_{\omega_{1}}]$, by virtue of an argument similar to that in the proof of Proposition 7.1, more details could be found in the proof of [25, Theorem 11]. ∎
@@ 8. Comments and open problems
Let ${\operatorname{P}(\omega)}$ be the power set of $\omega$, the set of natural numbers. We identify each element of ${\operatorname{P}(\omega)}$ with its characteristic function, an element of ${2^{\omega}}$. In that way we introduce a topology on ${\operatorname{P}(\omega)}$. Let ${[\omega]^{\omega}}$ be the family of all infinite sets in ${\operatorname{P}(\omega)}$. Each set in ${[\omega]^{\omega}}$ we identify with an increasing enumeration of its elements, a function in the Baire space ${\omega^{\omega}}$. We have ${[\omega]^{\omega}}\subseteq{\omega^{\omega}}$ and topologies in ${[\omega]^{\omega}}$ induced from ${\operatorname{P}(\omega)}$ and ${\omega^{\omega}}$ are the same. Let $\mathrm{Fin}$ be the family of all finite sets in ${\operatorname{P}(\omega)}$. A totally imperfect Menger set constructed by Bartoszyński and Tsaban, mentioned in Theorem 1.1 is a set the form $X\cup\mathrm{Fin}\subseteq{\operatorname{P}(\omega)}$ of size $\mathfrak{d}$, such that for any function $d\in{[\omega]^{\omega}}$, we have |$\left|\{\,x\in X:x\leq^{*}d\,\}\right|<\mathfrak{d}$. In fact, any set with these properties is totally imperfect and Menger [3, Remark 18.]. It has also a stronger covering property ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$, described in details in Subsection 8.2.
@@
Let $X\subseteq{[\omega]^{\omega}}$ be a set of size $\mathfrak{d}$ such that all coordinates of elements in $X$ are even numbers and for any function $d\in{[\omega]^{\omega}}$, we have |$\left|\{\,x\in X:x\leq^{*}d\,\}\right|<\mathfrak{d}$. Let $\{\,x_{\alpha}:\alpha<\mathfrak{d}\,\}$ be a bijective enumeration of elements in $X$. Fix a function $y\colon\mathfrak{d}\to{2^{\omega}}$. For each ordinal number $\alpha<\mathfrak{d}$, let $x_{\alpha}+y_{\alpha}$ be the function such that $(x_{\alpha}+y_{\alpha})(n):=x_{\alpha}(n)+y_{\alpha}(n)$ for all $n$. Then $x_{\alpha}+y_{\alpha}\in{[\omega]^{\omega}}$ and $x_{\alpha}\leq^{*}x_{\alpha}+y_{\alpha}$ for all ordinal numbers $\alpha<\mathfrak{d}$. Thus, for any function $d\in{[\omega]^{\omega}}$, we have $\{\,\alpha:x_{\alpha}+y_{\alpha}\leq^{*}d\,\}\subseteq\{\,\alpha:x_{\alpha}\leq^{*}d\,\}$, and the latter set has size smaller than $\mathfrak{d}$. Then the set $\{\,x_{\alpha}+y_{\alpha}:\alpha<\mathfrak{d}\,\}\cup\mathrm{Fin}$ is totally imperfect and Menger. It follows that for different functions $y\colon\mathfrak{d}\to{2^{\omega}}$, we get different totally imperfect Menger sets. ∎
@@
A cover of a space is a $\gamma$-cover if it is infinite and any point of the space belongs to all but finitely many sets in the cover. A space satisfies property ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of the space there are sets $U_{0}\in\mathcal{U}_{0},U_{1}\in\mathcal{U}_{1},\dotsc$ such that the family $\{\,U_{n}:n\in\omega\,\}$ covers the space. This property implies the Menger property. By the result of Just, Miller, Scheepers and Szeptycki [16, Theorem 2.3.] (see also the work of Sakai [23, Lemma 2.1.]), any subset of ${2^{\omega}}$ satisfying ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ is totally imperfect. The following questions are one of the major open problems in the combinatorial covering properties theory.

### v3

@@ 6. The weak grouped Menger game and Sacks forcing
For the proof of Theorem 6.1, we need the following auxiliary result. Similarly to Lemma 4.2, it can be proved in the same way as [6, Lemma 5.10], a rather standard argument is left to the reader.
@@
[…]
@@
Take $p$ and $F_{n},k_{n},l_{n},y_{\sigma_{n}}$ from Lemma 3.7, applied to $p_{0}$ and $\tau$. Let $\Sigma_{n}$ be the set of all maps $\sigma\colon F_{n}\to 2^{k_{n}}$ consistent with $p$, where $n\in\omega$. By Lemma 3.9(1), for each map $\sigma\in\Sigma_{n}$, there are maps $\sigma^{\prime}\neq\sigma^{\prime\prime}\in\Sigma_{n+1}$ extending $\sigma$, which implies $\sigma^{\prime}]\cap[\sigma^{\prime\prime}]=\emptyset$. For $S:=\operatorname{supp}(p)$, define a perfect set $K$, exactly in the same way as before Lemma 2.1, i.e.,
@@
Then the family of all sets $\sigma]$, where $\sigma\in\bigcup_{n\in\omega}\Sigma_{n}$, is a basis for $K$. Note also that Lemma 3.9(3) ensures that $S$ and the sequence $\langle\,k_{n},F_{n},\Sigma_{n}:n\in\omega\,\rangle$} satisfy $(e_{f})$ stated at the beginning of Section 2.
@@
Again, take $p$ and $F_{n},k_{n},l_{n},y_{\sigma_{n}}$ from Lemma 3.7, applied to $p_{0}$ and $\tau$. Let $\Sigma_{n}$, $\sigma]$ for $\sigma\in\Sigma_{n}$, $K$, and $h:K\to 2^{\omega}$ be defined in the same way as in the proof of (6.2.2). We have $p\Vdash\tau\in h[K].$
@@
The first half of Corollary 6.3(2), namely that all perfectly meager subsets of $2^{\omega}$} have cardinality at most $\omega_{1}$ in the Sacks model, was established by Miller [16, § 5].
@@
In $V[G_{\omega_{2}}]$, considered in the above corollary, there is a Luzin subset of $2^{\omega}$}, i.e., an uncountable set whose intersection with any meager set is at most countable, which is totally imperfect and Menger [13], but it is not (perfectly) meager. There is also a perfectly meager set that is not Menger: Since $\mathfrak{d}=\omega_{1}$ in $V[G_{\omega_{2}}]$, there is a dominating set $X=\{\,x_{\alpha}:\alpha<\omega_{1}\,\}$ in $\omega]^{\omega}$}, where $x_{\beta}\leq^{*}x_{\alpha}$ for all ordinal numbers $\beta<\alpha<\omega_{1}$. This set is not Menger. The set $X\cup\mathrm{Fin}$ satisfies the Hurewicz covering property [2, Theorem 10.], and any totally imperfect set with this property is perfectly meager [15, Theorem 5.5.]. Thus, the set $X$ is perfectly meager, too. Alternatively, we could use here the main result of [18], which implies directly that $X$ is perfectly meager since $\leq^{*}$ is a Borel subset of $\omega^{\omega}\times\omega^{\omega}$.
Theorems 4.1 and 6.1 motivate the following problem.
Problem 6.5.
In the Sacks model:
  1. (1)
  2. (2)
  3. (3)
    Are there only $\omega_{2}$-many Menger subsets of $2^{\omega}$?
  4. (4)
Regarding the last item of Problem 6.5, we do not know the answer even to the following question.
Problem 6.6.
@@ 7. Menger sets and Hechler forcing
The results from the previous section lead to the question, whether $\mathfrak{d}<\mathfrak{c}$ implies that any totally imperfect Menger subset of $2^{\omega}$} has cardinality at most $\mathfrak{d}$. We address this problem, showing that this is not the case. We also provide a consistent result that the size of the family of all Hurewicz subsets of $2^{\omega}$} can be equal $2^{\mathfrak{c}}$ even if $\mathfrak{d}<\mathfrak{c}$. Let $\mathbb{P}$ be a definable ccc forcing notion of size $\mathfrak{c}$ which adds dominating reals over a ground model (e.g., the poset defined in [3, p. 95], nowadays commonly named Hechler forcing) and $\mathbb{P}_{\omega_{1}}$ be an iterated forcing of length $\omega_{1}$ with finite support, where each iterand is equal to $\mathbb{P}$. In this section, by $G_{\omega_{1}}$ we mean a $\mathbb{P}_{\omega_{1}}$-generic filter over a ground model $V$ of ZFC.
@@
Let $X\subseteq{2^{\omega}}\cap V$ and $\phi\colon X\to{[\omega]^{\omega}}$ be a continuous function. The function $\phi$ can be extended to a continuous function $\Phi$ defined on a $G_{\delta}$-set $A$, containing $X$. Then there is an ordinal number $\alpha<\omega_{1}$ such that the function $\Phi$ and the set $A$ are coded in $V[G_{\alpha}]$. It follows that $\Phi[X]\subseteq V[G_{\alpha}]$. Since there exists a function $g_{\alpha}$, added in step $\alpha$, which dominates any real from $V[G_{\alpha}]$, the set $\Phi[X]$ is bounded in $V[G_{\omega_{1}}]$. By the Hurewicz–Recław characterization of the Hurewicz property [19, Proposition 1], the set $X$ is Hurewicz [^5] As noted by the referee, a similar argument actually gives that $\phi[X]$ is bounded for any Borel function $f:X\to\omega^{\omega}$. in $V[G_{\omega_{1}}]$. ∎
@@
[…]
@@
Since Cohen reals are added by any tail of the considered iteration, any subset of $2^{\omega}\cap V$ is also Rothberger in $V[G_{\omega_{1}}]$, by virtue of an argument similar to that in the proof of Proposition 7.1, more details could be found in the proof of [23, Theorem 11]. ∎
@@ 8. Comments and open problems
Let $\operatorname{P}(\omega)$} be the power set of $\omega$, the set of natural numbers. We identify each element of $\operatorname{P}(\omega)$} with its characteristic function, an element of $2^{\omega}$}. In that way we introduce a topology on $\operatorname{P}(\omega)$}. Let $\omega]^{\omega}$} be the family of all infinite sets in $\operatorname{P}(\omega)$}. Each set in $\omega]^{\omega}$} we identify with an increasing enumeration of its elements, a function in the Baire space $\omega^{\omega}$}. We have ${[\omega]^{\omega}}\subseteq{\omega^{\omega}}$ and topologies in $\omega]^{\omega}$} induced from $\operatorname{P}(\omega)$} and $\omega^{\omega}$} are the same. Let $\mathrm{Fin}$ be the family of all finite sets in $\operatorname{P}(\omega)$}. A totally imperfect Menger set constructed by Bartoszyński and Tsaban, mentioned in Theorem 1.1 is a set the form $X\cup\mathrm{Fin}\subseteq{\operatorname{P}(\omega)}$ of size $\mathfrak{d}$, such that for any function $d\in{[\omega]^{\omega}}$, we have $\left|\{\,x\in X:x\leq^{*}d\,\}\right|<\mathfrak{d}$. In fact, any set with these properties is totally imperfect and Menger [2, Remark 18.]. It has also a stronger covering property $\mathsf{S}_{1}(\Gamma,\mathrm{O})$}, described in details in Subsection 8.2.
@@
Let $X\subseteq{[\omega]^{\omega}}$ be a set of size $\mathfrak{d}$ such that all coordinates of elements in $X$ are even numbers and for any function $d\in{[\omega]^{\omega}}$, we have $\left|\{\,x\in X:x\leq^{*}d\,\}\right|<\mathfrak{d}$. Let $\{\,x_{\alpha}:\alpha<\mathfrak{d}\,\}$ be a bijective enumeration of elements in $X$. Fix a function $y\colon\mathfrak{d}\to{2^{\omega}}$. For each ordinal number $\alpha<\mathfrak{d}$, let $x_{\alpha}+y_{\alpha}$ be the function such that $(x_{\alpha}+y_{\alpha})(n):=x_{\alpha}(n)+y_{\alpha}(n)$ for all $n$. Then $x_{\alpha}+y_{\alpha}\in{[\omega]^{\omega}}$ and $x_{\alpha}\leq^{*}x_{\alpha}+y_{\alpha}$ for all ordinal numbers $\alpha<\mathfrak{d}$. Thus, for any function $d\in{[\omega]^{\omega}}$, we have $\{\,\alpha:x_{\alpha}+y_{\alpha}\leq^{*}d\,\}\subseteq\{\,\alpha:x_{\alpha}\leq^{*}d\,\}$, and the latter set has size smaller than $\mathfrak{d}$. Then the set $\{\,x_{\alpha}+y_{\alpha}:\alpha<\mathfrak{d}\,\}\cup\mathrm{Fin}$ is totally imperfect and Menger. It follows that for different functions $y\colon\mathfrak{d}\to{2^{\omega}}$, we get different totally imperfect Menger sets. ∎
@@
A cover of a space is a $\gamma$-cover if it is infinite and any point of the space belongs to all but finitely many sets in the cover. A space satisfies property $\mathsf{S}_{1}(\Gamma,\mathrm{O})$} if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of the space there are sets $U_{0}\in\mathcal{U}_{0},U_{1}\in\mathcal{U}_{1},\dotsc$ such that the family $\{\,U_{n}:n\in\omega\,\}$ covers the space. This property implies the Menger property. By the result of Just, Miller, Scheepers and Szeptycki [15, Theorem 2.3.] (see also the work of Sakai [21, Lemma 2.1.]), any subset of $2^{\omega}$} satisfying $\mathsf{S}_{1}(\Gamma,\mathrm{O})$} is totally imperfect. The following questions are one of the major open problems in the combinatorial covering properties theory.